OpenStudy (joftheworld):

@ridoleor What is the equation (in slope-intercept form) of a line that passes through the point (3, -2) and has a slope of -2?

OpenStudy (anonymous):

y + 2 = -2(x-3) y + 2 = -2x + 6 y = -2x + 4

OpenStudy (joftheworld):

thanks, what about.. What is the equation of a line that passes through (-6, 2) and has a slope of -1/2

OpenStudy (anonymous):

y - 2 = -1/2(x +6) y - 2 = -1/2x - 3 y = -1/2x - 1

OpenStudy (anonymous):

Automatically you can tell it's the last one, because that's the only one with a positive slope.

OpenStudy (joftheworld):

sorry im not very good with graphs..

OpenStudy (anonymous):

It's alright. For this one, you can eliminate the last 2 because they are not in point-slope form anyway. Then you can eliminate the first one, because it has a positive slope.

OpenStudy (joftheworld):

thanks, and

OpenStudy (anonymous):

The second one; it's the only one with a neg. slope

OpenStudy (joftheworld):

What is the equation of a line that passes through (4, 7) and has a slope of -1/4